OpenStudy (anonymous):

HELP im lost!!! What is the radius of a circle given by the equation (x+1)^2+(y-2)^2=64

OpenStudy (amistre64):

whats the general equation of a circle?

OpenStudy (precal):

radius is 8

OpenStudy (anonymous):

is it? how?

OpenStudy (amistre64):

good question :)

OpenStudy (amistre64):

the general equation of a circle is defined as: x^2+y^2 = r^2 so that that last bit is the radius

OpenStudy (amistre64):

if we compare this to what you have; that amounts to r^2 = 64; therefore r = 8

OpenStudy (anonymous):

yea so do you square the radius at the end? just that?

OpenStudy (amistre64):

yep; other equation might have to be worked INTO this form, but if they are already looking like this then the radius is the sqrt of the last bit

OpenStudy (anonymous):

ooohhh ok

OpenStudy (anonymous):

thank you i got it!:D

OpenStudy (amistre64):

good ;)